OpenStudy (anonymous):

Will Fan And Medal Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of f[g(5)]?

OpenStudy (anonymous):

312 70 168 −1

OpenStudy (decentnabeel):

f(x) = 3x − 2 and g(x) = x2 − 1 first put x=5 in g(x) g(x)=x^2-1 g(5)=5^2-1 g(5)=25-1 =24

OpenStudy (decentnabeel):

then f(x)=3x-2 put x=g(5) f(g(5)=3(24)-2 =72-2 =70 that is the answer

OpenStudy (decentnabeel):

@fratom123 are you understand